# spacing=nonfrench

### v3

@@ Abstract.
5G mmWave roadside networks promise high bandwidth but face significant challenges in maintaining reliable connections for users moving at high speed. Frequent handovers, complex beam alignment, and signal blockage from car bodies lead to service interruptions and degraded performance. We present Wall-Street, a vehicle-mounted smart surface that enhances 5G mmWave connectivity for in-vehicle users. Wall-Street improves mobility management by (1) steering outdoor mmWave signals into the vehicle for shared coverage, enabling a single, collective handover decision for all users; (2) enabling neighbor-cell search without interrupting data transfer, allowing for seamless handovers; and (3) connecting users to a new cell before disconnecting from the old cell for reliable handovers. We have implemented and integrated Wall-Street in the COSMOS testbed and evaluated its real-time performance with multiple gNBs and users inside a surface-mounted vehicle, driving on a nearby road. In multi-UE scenarios, Wall-Street improves TCP throughput by up to 78% and reduces RTT by up to 34% over the standard 5G Standalone handover protocol.
@@ 1. Introduction
5G wireless networks leverages millimeter wave (mmWave) spectrum in the FR2 band to deliver multi-gigabit, low-latency connectivity (3GPP, 2022). However, due to its short wavelengths and limited diffraction capability, mmWave spectrum suffers from high path loss and extreme vulnerability to blockage, posing significant challenges in maintaining reliable connections, particularly in high mobility scenarios (Rappaport et al., 2017).
To overcome poor signal propagation, mmWave networks require dense deployments of small cells, which in turn leads to handovers (HOs) that are far more frequent than in low-band networks. Indeed, mmWave 5G NR experiences a HO significantly more often than LTE or low band 5G NR (Hassan et al., 2022; Narayanan et al., 2021). The mmWave HO process adds a layer of complexity on top of low-band procedures (Deng et al., 2018; Xu et al., 2020), requiring precise alignment of narrow beams between the base station (gNB) and user equipment (UE) (Tassi et al., 2017). To achieve this, the standard HO protocol requires the UE to frequently suspend its data connection to measure neighboring gNBs, a procedure that directly leads to service interruptions and performance degradation.
High delays associated with mmWave HOs have significant impacts throughout the entire protocol stack, from the physical layer to the application layer, degrading quality of service (QoS) for UEs, especially for latency-sensitive applications (Hassan et al., 2022; Narayanan et al., 2020). To prepare a HO, the UE immediately stops data communication with the serving gNB to perform neighboring cell measurements. This process requires UEs to engage in an exhaustive search for the strongest signal from neighboring gNBs, resulting in bufferbloat and slow growth of the congestion window (Mezzavilla et al., 2018). While existing works (Wang et al., 2020) primarily focus on reducing beam searching overhead, the fundamental problem of maintaining reliable connectivity remains unexplored. Also, the HO decision relies on these measurement from UEs, which may be outdated or lost due to the rapidly fluctuating channel conditions. This can trigger flawed decisions, leading to ping-pong handovers and connection timeouts, impacting overall network performance.
This paper presents the design and implementation of Wall-Street, a programmable smart surface strategically deployed on a vehicle’s interior to address these challenges. Wall-Street acts as an intelligent gateway for mmWave signals, overcoming vehicle body blockage and redesigning the HO process to eliminate service interruptions for in-vehicle users. It achieves this through a new capability called dual-beam control; Wall-Street steers two incoming beams simultaneously in reflective and/or transmissive directions, with arbitrary power division between them. Using this ability, it offloads the HO task from individual UEs to the surface itself. This allows all in-vehicle users to maintain data connectivity while the network performs HO process on their behalf. As illustrated in Fig.˜1, we summarize three key innovations:
@@
(b) Transflexive handover, ongoing comm.
(c) Enhance handover reliability.
Figure 1. Wall-Street’s design innovations electronically split, shape, and steer mmWave transmissions in real time to enable seamless roadside 5G networks.
1. Batched handover via shared in-vehicle coverage. Wall-Street provides consistent, shared mmWave coverage to all UEs by steering outdoor signals into the vehicle (Fig.˜1(a)). This also allows the network to batch the HO procedure for all in-vehicle UEs into a single operation, eliminating the overhead and contention of separate HOs.
2. Seamless handover with concurrent communication. Innovating beyond existing surfaces that can only reflect or refract at one time (Cho et al., 2023), Wall-Street uniquely steers two mmWave beams simultaneously in transmissive and reflective angles with minimal signal loss. Specifically, it reflects neighboring gNB’s beacon signals to the serving gNB for measurement while keeping the user data link through the transmitted path [^1] To avoid collisions, nearby gNBs separate their respective signals in frequency and time across different gNBs, allowing the source gNB to listen to other cells while simultaneously exchanging data with the UE. (Fig.˜1(b)). This eliminates the standard protocol’s requirement to suspend data transmission during cell search, removing service interruption. Since users are not involved in the scanning process, it also minimizes UE energy consumption related to HOs, which typically consumes twice the energy compared to low-band HOs (Hassan et al., 2022).
3. Make-before-break handover. At the moment of HO, Wall-Street combines beams from both the source and target gNBs. By steering duplicate data packets from both gNBs to the user, it significantly reduces packet loss and retransmissions, enhancing reliability during the transition. Also, by adjusting the power of the respective links, Wall-Street ensures an uninterrupted flow of data.
We have designed, implemented, and integrated the full Wall-Street system, including its hardware and protocol, into the COSMOS wireless testbed (Raychaudhuri et al., 2020). We have also implemented new PHY features on the testbed for HO operations. We mount Wall-Street on the rear door of an SUV with multiple mobile UEs inside, communicating with three gNBs placed on the first floor of a lab facing the road. To evaluate TCP end-to-end performance, we use a state-of-the-art mmWave ns-3 simulator  (Mezzavilla et al., 2018) with our own modifications. The simulation is trace-driven, fed with PHY traces collected from real-time experiments. We compare our system against a standard 5G Standalone HO protocol that we also implemented as a baseline. In multi-user scenarios, our results demonstrate that Wall-Street improves TCP throughput by up to 78%, reduces RTT by up to 34%, and cuts unnecessary HOs in half. Microbenchmarks validate Wall-Street’s in-vehicle coverage improvement, beam tracking, and dual-beam steering.
Contribution. Our main contributions are as follows:
@@
    A dual-link beam manipulation. We design, build, and demonstrate the first mmWave smart surface capable of simultaneously and independently steering a transmissive link (into the vehicle) and a reflective link (between gNBs) with programmable power division.
@@
    Reliable make-before-break handover. We realize a make-before-break HO protocol that leverages our surface to connect UEs to a new gNB before disconnecting from the old gNB. By receiving packets from both gNBs, our system improves HO reliability.
@@
    Scalable, vehicle-centric handover. We introduce an algorithm that makes a single, batched HO decision for all in-vehicle users, reducing signaling overhead. By estimating the stable gNB-to-surface link quality from multi-user measurements, our design also improves HO decision.
@@
    Real-world system implementation. We implement the full Wall-Street hardware system and integrate it in the COSMOS testbed. Our experiments with a surface-mounted SUV and multi-gNBs/UEs reveals significant improvement over the standard protocol.
@@ 2. Related Work
High-speed Mobility Management. Several works address sub-6 GHz vehicular mobility. Wi-Fi Goes to Town (Song et al., 2017) provides continuous connectivity via rapid switching between roadside Wi-Fi APs. POLYCORN (Ni et al., 2023) optimizes traffic scheduling for high-speed railways but requires fixed trajectories, making it unsuitable for unpredictable vehicular movement. REM (Li et al., 2020) uses multipath profiles instead of signal strength for HO decisions. However, these systems target sub-6 GHz bands, where signals penetrate obstacles easily with minimal cell search overhead. At mmWave, car bodies block signals completely and narrow beams lose alignment easily, demanding fundamentally different solutions
@@
[…]
Mobile mmWave Repeaters. Commercial mmWave repeaters for homes (Commware, 2020) and vehicles (Hyatt, [n. d.]) use external antennas to receive signals, amplify, and retransmit them to users inside. However, they consume 20-40 W, far exceeding Wall-Street’s 160 µW max power consumption [^2] With COTS DAC controllers (noa, [n. d.]), total power consumption ranges from 0.48 to 0.64 W. Custom DAC designs may further reduce the power consumption.. They cannot support concurrent transmission/reflection or make-before-break handover since they cannot relay two links at the same time. Such operation requires additional power-hungry RF chains. Moreover, repeaters requires higher search complexity compared to Wall-Street (Section˜4.2.1). Alternatively, rooftop-to-WiFi relay could bottleneck mmWave’s multi-Gbps to WiFi’s data rate while adding offloading latency.
@@
Reconfigurable Intelligent Surface (RIS). RISs manipulate mmWave signals through programmable metamaterials that directly impart phase shifts on incoming waves rather than generating new signals, consuming minimal power. As seen in Table˜1, (Qian et al., 2022) is a passive RIS, reflecting signals to fixed directions for blind spots, unable to support mobile UEs. Other reflective RISs steer signals through switches (Tan et al., 2018), PIN diodes (Yang et al., 2022; Gros et al., 2021, 2023), or varactor diodes  (Wolff et al., 2023; Kim et al., 2024; Cho et al., 2023). While (Kim et al., 2024) integrates 5G protocols, it only reflects, not suitable for in-vehicle use. (Cho et al., 2023) is the first mmWave surface that can steer in both refraction and reflection but requires time-multiplexing between modes. Wall-Street supports simultaneous operation with arbitrary power division.
@@
Roadside RIS deployments  (Chen et al., 2021; Alsenwi et al., 2022; Tian et al., 2022) could help gNBs overcome environmental blockage but require extensive infrastructure and cannot solve vehicle body penetration. Wall-Street, mounted directly on vehicles, moves with users to ensure continuous in-vehicle coverage and reliable HOs.
3. Primer: 5G SA Handover
Figure 2. 5G SA Xn handover: (1) Preparation phase; (2) Execution phase; (3) Completion phase. Data exchange is denoted with orange and yellow rectangles.
@@ 3. Primer: 5G SA Handover
5G Standalone (SA) uses Xn handover, where gNBs exchange information related to the HO process via the Xn interface. The process has three phases as outlined in Fig.˜2: Preparation, Execution, and Completion.
(1) Preparation phase. 5G SA relies on UE-side feedback to trigger handovers. The serving cell periodically requests UEs to stop ongoing transmission and measure Reference Signal Received Power (RSRP) from neighboring cells. UEs search exhaustively, then send measurement reports (MR) containing neighbor IDs and RSRPs back to the serving gNB. This process disrupts service and drains UE battery.
A common trigger for 5G HO is event A3 (Deng et al., 2018; Xu et al., 2020). It triggers when neighbor cell RSRP ($M_{n}$) exceeds serving cell RSRP ($M_{s}$) by a certain hysteresis threshold $H$: $M_{n}>M_{s}+H$. To prevent ping-pong HO from signal fluctuations, HO only executes when A3 holds true for the time-to-trigger (TTT) duration.
@@
(2) Execution phase. The serving gNB sends a handover request to the chosen target gNB via Xn interface. Upon acknowledgment, the serving gNB commands the UE to detach from itself and synchronize with the target using Random Access (RACH), while forwarding buffered data to the target gNB. However, handovers typically occur at edges where both links are weak, causing packet loss and retransmissions.
@@
(3) Completion phase. The target gNB requests path switching to the 5G Core, then signals the source gNB to release the UE context via Xn interface. Poor HO decisions lead to ping-pong handovers or TCP connection termination.
@@
4. Design
This section presents the design of Wall-Street, including the surface design (Section˜4.1) and HO system design (Section˜4.2). The serving gNB controls Wall-Street via sub-6 GHz control channel.
4.1. Wall-Street Hardware
Surface Design. Wall-Street is a 26 GHz Huygens metasurface (HMS) that supports both reflection and transmission. HMS consists of orthogonal electric and magnetic meta-atoms facing each other across a dielectric substrate (shown in Fig.˜3 on left), creating discontinuities in EM fields (Cho et al., 2023) [^3] We refer the reader to Appendix A for the details on our design parameters. Demonstration videos are available in bit.ly/42xb2nr. This structure can manipulate the magnitude and phase of the incident field in both reflective and transmissive directions. Each meta-atom contains a low-power voltage-controlled capacitor. By adjusting the bias voltages applied to the magnetic and electric meta-atoms $(u_{m},u_{e})$, we can change the complex reflection and transmission coefficients of the incoming signal. Figure˜3 (right) shows Vector Network Analyzer (VNA) measurements of Wall-Street’s response to different voltage combinations. The marked regions (black arrows) indicate where we achieve arbitrary phase shifts from 0 to $2\pi$ while maintaining high efficiency in reflection or transmission.
Dual-Link Beam Steering. Wall-Street’s key innovation is transflective operation; it simultaneously steers one beam in transmission (refraction) and another in reflection while adjustable power division between them. The challenge is to find an optimal configuration that jointly satisfies these multiple objectives, achieving two steering angles with a desired power ratio.
A straightforward approach is to spatially partition the surface into two - for example, assigning 80% of the area to the transmissive link and 20% to the reflective link. A distinct phase gradient is then applied to each partition to steer the two beams independently. While it is computationally fast, hard partitioning raises undesirable sidelobes and, based on our own experimental tests, reduces gain for both links.
Instead of hard partitioning, we use a soft partitioning approach that optimizes the contribution of individual meta-atom to both links. We use an optimization algorithm to find a single voltage configuration for the entire surface that maximizes the overall gain in the two desired directions while maintaining the power ratio and suppressing sidelobes. These configurations are pre-computed offline and stored in a codebook for real-time operation.
To populate this codebook, we find the optimal voltages $\left\{(u_{m,n},u_{e,n})\right\}_{n=0}^{N-1}$ for a desired pair of steering angles ($\theta_{1}$, $\theta_{2}$) with a power split ratio $\alpha$ using a genetic algorithm:
@@ 4.1. Wall-Street Hardware
where $\phi_{k}$ is the ideal phase shift required to steer a beam by $\theta_{k}$, and $c_{t,n}$, $c_{r,n}$ are the complex transmission and reflection coefficients of $n$-th meta-atom from S-parameter measurements (Fig.˜3). The resulting codebook $\Theta$, indexed by $(\theta_{t},\theta_{r},\alpha)$, is used during real-time operation (Section˜4.2). For dual-transmission, the reflective term is simply replaced with a second transmissive term, and for single-beam steering, one term is removed entirely.
@@
Wall-Street’s dual-link beam steering, which independently controls two angles across a wide $\pm 70\degree$ range, provides two functions in our handover protocol, as seen in Fig.˜4. During neighbor cell search (Fig.˜4(a)), Wall-Street maintains a transmissive link between the serving gNB and UE (blue) while reflecting neighbor cell’s broadcast signals back toward the serving gNB (red) for cell search. During a make-before-break HO (Fig.˜4(b)), Wall-Street steers both the serving and target gNB’s signals to the UE. Angular reciprocity, where the surface maintains identical beam paths for uplink and downlink, enables instantaneous uplink/downlink switching without reconfiguration. In contrast, repeaters must switch between different receive and transmit beam patterns, adding latency.

### v4

@@ Abstract.
mmWave networks promise high bandwidth but face significant challenges in maintaining reliable connections for users moving at high speed. Frequent handovers, complex beam alignment, and signal blockage from car bodies lead to service interruptions and degraded performance. We present Wall-Street, a vehicle-mounted smart surface that enhances mmWave connectivity for in-vehicle users. Wall-Street improves mobility management by (1) steering outdoor mmWave signals into the vehicle for shared coverage and providing a single, collective handover for all users; (2) performing neighbor-cell search without interrupting data transfer, ensuring seamless handovers; and (3) connecting users to a new cell before disconnecting from the old cell for reliable cell transitions. We implemented and integrated Wall-Street into the COSMOS testbed. We collected PHY traces with multiple base station nodes and in-vehicle user nodes with a surface-mounted vehicle, driving on a nearby road. Our trace-driven ns-3 simulation demonstrates a throughput improvement of up to 78% and a latency reduction of up to 34% over the standard Standalone handover scheme. [^1] This work was done when Kun Woo Cho was a student at Princeton University.
@@ 1. Introduction
5G wireless networks leverages millimeter wave (mmWave) spectrum in the FR2 band to deliver multi-gigabit, low-latency connectivity (3GPP, 2022b). However, due to its short wavelengths and limited diffraction capability, mmWave spectrum suffers from high path loss and extreme vulnerability to blockage, posing significant challenges in maintaining reliable connections, particularly in high mobility scenarios (Rappaport et al., 2017).
To overcome poor signal propagation, mmWave networks require dense deployments of small cells, which in turn leads to handovers (HOs) that are far more frequent than in low-band networks. Indeed, mmWave 5G NR experiences a HO significantly more often than LTE or low band 5G NR (Hassan et al., 2022; Narayanan et al., 2021). The mmWave HO process adds a layer of complexity on top of low-band procedures (Deng et al., 2018; Xu et al., 2020), requiring precise alignment of narrow beams between the base station (gNB) and user equipment (UE) (Tassi et al., 2017).
High delays associated with mmWave HOs have significant impacts throughout the entire protocol stack degrading quality of service (Hassan et al., 2022; Narayanan et al., 2020). To determine when a handover is needed, the standard protocol requires users to pause data communication, scan neighboring cells, and report the results to the serving cell. This process requires users to engage in an exhaustive search for the strongest signal from neighboring cells, resulting in service interruptions and performance degradation (Mezzavilla et al., 2018). While existing works (Wang et al., 2020) primarily focus on reducing beam searching overhead, the fundamental problem of maintaining reliable connectivity remains unexplored. Also, the HO decision relies on these measurement from users, which may be outdated or lost due to the rapidly fluctuating channel conditions. This can trigger flawed decisions, leading to ping-pong handovers and connection timeouts, impacting overall network performance.
This paper presents the design and implementation of Wall-Street, a programmable smart surface strategically deployed on a vehicle’s interior to address these challenges. Wall-Street acts as a programmable L1 extension of base stations that overcomes vehicle body blockage and makes the HO process transparent to users. It achieves this through a new PHY-layer capability called dual-beam control. Wall-Street steers two incoming beams simultaneously in reflective and/or transmissive directions, with arbitrary power division between them. Using this ability, it offloads the HO task from individual user to the surface itself. This allows all in-vehicle users to maintain data connectivity while the network performs HO process on their behalf. As illustrated in Fig.˜1, we summarize three key innovations:
@@
(b) Concurrent communications and cell scanning.
(c) Make-before-break handover.
Figure 1. Wall-Street’s design innovations electronically split, shape, and steer mmWave transmissions in real time to ensure seamless roadside mmWave networks.
1. Batched handover via shared in-vehicle coverage. Wall-Street provides consistent, shared mmWave coverage to all users by steering outdoor signals into the vehicle (Fig.˜1(a)). This also allows the network to batch the HO procedure for all in-vehicle users into a single operation, eliminating the overhead and contention of separate HOs.
2. Seamless handover with concurrent communication. Innovating beyond existing surfaces that can only reflect or refract at one time (Cho et al., 2023), Wall-Street uniquely steers two mmWave beams simultaneously in transmissive and reflective angles. It reflects neighboring cell’s broadcast signals to the serving cell while keeping the user data link through a transmissive path (Fig.˜1(b)). This allows the serving cell to measure neighbor cells directly. By offloading the scanning burden from the user, our system eliminates service interruptions and significantly reduces UE power consumption associated to HOs, which consumes twice the energy compared to low-band HOs (Hassan et al., 2022).
3. Make-before-break handover. At the moment of HO, Wall-Street maintains links from both the old and new cells. Users receive duplicate packets from both cells, reducing packet loss and retransmissions and enhancing reliability during cell transitions [^2] To avoid collisions, nearby cells separate their respective signals in frequency and time. This allows the serving cell to listen to other cells during data communication and the user to receive signals from two cells..
We have designed, implemented, and integrated the Wall-Street system into the COSMOS wireless testbed (Raychaudhuri et al., 2020). We have also implemented new PHY features on the testbed. We mount Wall-Street on the rear door of an SUV with multiple mobile users inside, communicating with three cells placed on the first floor of a lab facing the road. To evaluate TCP performance, we use a mmWave ns-3 simulator (Polese, 2019). We modified the simulator to be trace-driven, fed with PHY traces collected from driving experiments. In multi-user scenarios, our ns-3 results demonstrate that Wall-Street improves TCP throughput by up to 78%, reduces RTT by up to 34%, and cuts unnecessary HOs in half. Our microbenchmarks validate Wall-Street’s in-vehicle coverage improvement, beam tracking, and dual-beam steering [^3] GitHub codes: https://github.com/kunwooch/Wall-Street [^4] Demonstration videos: https://youtu.be/35iiBFYVkZY.
Contributions. Our main contributions are as follows:
@@
    A dual-link beam manipulation. We design, build, and demonstrate the first mmWave smart surface capable of simultaneously and independently steering a transmissive link (into the vehicle) and a reflective link (between cells) with programmable power division.
@@
    Reliable make-before-break handover. We realize a make-before-break handover via Wall-Street, which connects in-vehicle users with two cells. By receiving packets from both cells, our system improves reliability.
@@
    Scalable, vehicle-centric handover. We introduce an algorithm that makes a single, batched handover for all in-vehicle users, reducing signaling overhead. We leverage measurements from multiple users to estimate the cell-to-surface link quality and make a reliable handover decision.
@@
    Real-world system implementation. We implemented the Wall-Street hardware and integrated it within the COSMOS testbed. We conducted driving experiments with a surface-equipped SUV and multiple gNB/UE nodes.
@@ 2. Related Work
High-speed Mobility Management. Several works address sub-6 GHz vehicular mobility. Wi-Fi Goes to Town (Song et al., 2017) provides continuous connectivity via rapid switching between roadside Wi-Fi APs. POLYCORN (Ni et al., 2023) optimizes traffic scheduling for high-speed railways but requires fixed trajectories, making it unsuitable for unpredictable vehicular movement. REM (Li et al., 2020) uses multipath profiles instead of signal strength for HO decisions. However, these systems target sub-6 GHz bands, where signals penetrate obstacles easily with minimal cell search overhead. At mmWave, car bodies block signals completely and narrow beams lose alignment easily, demanding fundamentally different solutions.
[…]
@@
Mobile mmWave Repeaters. Commercial mmWave repeaters for homes (Commware, 2020) and vehicles (Hyatt, 2021) use external antennas to receive signals, amplify, and retransmit them to users inside. This architecture relies on digital processing and active RF chains, resulting in a power consumption of 20–40 W. This far exceeds Wall-Street’s 160 µW power [^5] With two COTS DACs, the total power consumption is 0.48 W. We refer to Sec. 7 for a detailed power analysis.. Also, controlling two mmWave beams simultaneously would require repeaters to add more RF chains, further increasing power. Repeaters also incur higher search complexity (Section˜4.2.1). Finally, rooftop-to-WiFi relays are suboptimal as they introduce latency and limit mmWave’s multi-Gbps throughput to WiFi speeds.
@@
Reconfigurable Intelligent Surface (RIS). RISs controls mmWave signals using programmable metamaterials that directly phase-shift incoming waves. As seen in Table˜1, (Qian et al., 2022) uses passive RISs to reflect signals in fixed directions but unable to support mobile users. More advanced reflective RISs use switches (Tan et al., 2018), PIN diodes (Yang et al., 2022; Gros et al., 2021, 2023), or varactor diodes  (Wolff et al., 2023; Kim et al., 2024; Cho et al., 2023) for beam steering. While (Kim et al., 2024) integrates 5G protocols, they are limited to reflection mode, making them unsuitable for in-vehicle coverage. (Cho et al., 2023) introduces the ability to switch between transmission and reflection but requires time-multiplexing. Wall-Street supports simultaneous transmission and reflection with flexible power division.
@@
While roadside RIS deployments (Chen et al., 2021; Alsenwi et al., 2022; Tian et al., 2022) help mitigate environmental blockage but require extensive infrastructure and cannot solve vehicle body penetration. Wall-Street, mounted directly on vehicles, moves with users to ensure continuous in-vehicle coverage and reliable handovers.
3. Primer: Standalone Handover
@@
Standalone (SA) base stations exchange handover signaling via the Xn interface. The process has three phases: Preparation, Execution, and Completion.
(1) Preparation phase. The network relies on UE measurements to trigger handovers. The serving cell periodically requests UEs to pause ongoing transmissions to measure the Reference Signal Received Power (RSRP) of neighboring gNBs. The UE performs an exhaustive search and sends a Measurement Report (MR) containing neighboring gNB IDs and RSRP values back to the serving gNB. A handover is typically triggered by event A3 (Deng et al., 2018; Xu et al., 2020), which occurs when the neighboring gNB’s RSRP ($M_{n}$) exceeds the serving cell’s RSRP ($M_{s}$) by a hysteresis threshold $H$ ($M_{n}>M_{s}+H$). To prevent ping-pong HO from signal fluctuations, HO executes when A3 holds true for the time-to-trigger (TTT) duration.
@@
(2) Execution phase. The serving gNB sends a handover request to the selected target gNB via the Xn interface. Upon acknowledgment, the serving gNB commands the UE to detach from itself and synchronize with the target using Random Access (RACH), while forwarding buffered data to the target gNB. Because this transition often occurs at the service boundary, the newly established link suffers from packet loss and retransmissions.
@@
(3) Completion phase. The target gNB requests path switching to the Core and signals the serving gNB to release the UE context via the Xn interface. Poor HO decisions often lead to ping-pong handovers or TCP connection termination.
4. Design
This section presents the design of Wall-Street, including the surface design (Section˜4.1) and handover design (Section˜4.2).
4.1. Wall-Street Hardware
Wall-Street has three key contributions: (1) Wall-Street operates at 26 GHz with dual-link steering and power division; (2) Wall-Street provides an adaptive control layer that is transparent to the user. Unlike conventional relays requiring active UE coordination, Wall-Street externalizes UE control, offloading mobility management to the infrastructure; (3) by manipulating links in the analog domain, Wall-Street minimizes the energy and latency of digital processing required by active relays (e.g., receiving, decoding, encoding, and re-transmitting).
Surface Design. Wall-Street is a 26 GHz Huygens metasurface (HMS) that supports both reflection and transmission with the gain of $29$ dB. HMS consists of orthogonal electric and magnetic meta-atoms facing each other across a dielectric substrate (shown in Fig.˜2 on left), creating discontinuities in EM fields (Cho et al., 2023). This structure can manipulate the magnitude and phase of the incident field in both reflective and transmissive directions. Each meta-atom contains a low-power voltage-controlled capacitor. By adjusting the bias voltages applied to the magnetic and electric meta-atoms $(u_{m},u_{e})$, we can change the complex reflection and transmission coefficients of the incoming signal. Figure˜2 (right) shows Vector Network Analyzer (VNA) measurements of Wall-Street’s response to different voltage combinations. The marked regions (black arrows) indicate where we achieve arbitrary phase shifts from 0 to $2\pi$ while maintaining high efficiency in reflection or transmission.
@@ 4.1. Wall-Street Hardware
Dual-Link Beam Steering. Wall-Street simultaneously steers one beam in transmission and another in reflection with adjustable power splitting. The challenge is to find an optimal configuration that jointly satisfies these multiple objectives, achieving two steering angles with a desired power ratio.
A straightforward approach is to spatially partition the surface into two - for example, assigning 80% of the area to the transmissive link and 20% to the reflective link. A distinct phase gradient is then applied to each partition to steer the two beams independently. While it is computationally fast, hard partitioning raises undesirable sidelobes and reduces gain for both links. Instead, we use soft partitioning, which optimizes the entire surface to serve both objectives simultaneously. This approach finds a single voltage configuration that maximizes gain in both directions while maintaining the desired power split and suppressing sidelobes.
We pre-compute these configurations using a genetic algorithm and store them in a lookup table (codebook) for real-time access. To populate this codebook, we find the optimal voltages $\left\{u_{m,n},u_{e,n}\right\}$ that maximize the combined array factor for a target pair of angles ($\theta_{1}$, $\theta_{2}$) and power split $\alpha$:
@@
where $\phi_{k}$ is the ideal phase shift for steering to angle $\theta_{k}$. The terms $c_{t,n}$ and $c_{r,n}$ denote the complex transmission and reflection coefficients of the $n$-th meta-atom, derived from S-parameter measurements in Fig.˜2 (right). The resulting codebook $\Theta$, indexed by $(\theta_{t},\theta_{r},\alpha)$, is used during real-time operation described in Section˜4.2. For dual-transmission, the reflective term is replaced by a second transmissive term. For single-beam steering, the second term is removed entirely.
@@
Wall-Street’s dual-link beam steering, which independently controls two angles across a wide $\pm 70\degree$ range, provides two handover functions, as illustrated in Fig.˜3: (1) neighbor cell search (Fig.˜3(a)): Wall-Street maintains a transmissive link between the serving gNB and UE (blue) while reflecting neighbor cell’s broadcast signals back toward the serving gNB (red) for cell search, and (2) make-before-break handover (Fig.˜3(b)): Wall-Street steers both the serving and target gNB’s signals to the UE. Furthermore, we leverage surface’s angular reciprocity, where the beam path is identical for uplink and downlink. This allows for instantaneous TDD switching without surface reconfiguration. In contrast, active repeaters must switch between receive and transmit beam patterns, introducing additional latency